# Sequential Monte Carlo Squared for Online Inference in Stochastic Epidemic Models

### v1

@@ 3.2.2 Data and model fitting
The data source used to make inferences was restricted to daily confirmed cases counts extracted from the updated database of the Health Protection Surveillance Centre (HPSC) (https://COVID19ireland-geohive.hub.arcgis.com/). To integrate these observations into our model, we define two variables: $C^{I}_{t}$ representing the cumulative number of infectious cases. These are given by the following equation:
@@
Hence, $\mu_{t}=C^{I}_{t+1}-C^{I}_{t}$ represents the daily expected number of cases. Since daily reported case counts are subject to observation errors and typically exhibit overdispersion (where the conditional variance exceeds the conditional mean), we model the observed daily cases, $Z^{I}_{t}$, using a normal approximation (chosen for computational efficiency, Funk et al. (2018)) of the negative binomial distribution. The negative binomial distribution is a generalization of the Poisson distribution that accounts for overdispersion in the observation. Specifically, for the negative binomial distribution, the mean is $\mu_{t}$ and the variance is $\mu_{t}(1+\phi\mu_{t})$, where $\phi$ is the overdispersion parameter. To avoid singularities when $\mu_{t}=0$, we followed the approach of Funk et al. (2018), rounding variances smaller than 1 up to 1. Thus, the observed daily case counts follow a negative binomial distribution:
@@
We estimate the spread of COVID-19 in Ireland from February 29th, 2020, to February 28th, 2021, using the daily new infected count data of 365 days. The Irish population in 2020 was estimated to be $N=4,965,439$. The O-SMC2 uses $N_{\theta}=1000$ parameter particles, $N_{x}=500$ state particles, $M=5$ successive PMMH moves and a window size of $t_{k}=80$ days. Table C.1 gives the description and prior of the inferred parameters used in the model (3.2.1). These priors are based on previous studies of COVID-19 dynamics in Ireland (Irish Department of Health, 2020; Cazelles et al., 2021). For the choice of prior distributions for the hyperparameters, $\nu_{\beta}$ (which governs the variability in the transmission rate) and $\phi$ (the overdispersion parameter), we adopt settings that allow for rapid changes in the transmission rate, ensuring the model can adapt to sudden shifts in disease dynamics, as discussed in (Funk et al., 2018). Regarding initial states, we set: $S_{0}=N-E_{0}-A_{0}$, $E_{0}=1$, $A_{0}\sim\mathcal{U}(\{10,\dots,50\})$ and $I_{0}=R_{0}=0$. The execution time for the 365 days considered here was less than 5 hours, making it suitable for practical real-time applications.
@@
Figure 5 illustrates the model’s fit to the daily number of COVID-19 cases reported in Ireland. The top plot shows the daily new infections, with three noticeable peaks: one in mid-April 2020, another in mid-November 2020, and a larger peak in early January 2021. The model successfully captures these three waves, with observed data (represented by the orange dots) falling within the 95% credible interval on most days. However, a sudden drop in reported cases around mid-January 2021 is not fully captured by the model. This anomaly could be due to the smoothing effect of Gaussian noise in the model, which limits its ability to promptly detect abrupt changes. Additionally, the drop may also reflect real-world factors such as reporting delays, testing changes, or a shift in public health measures that were not directly incorporated into the model. Beyond March 31, 2021, 2021, the model projects cases forward for 21 days, with data from April 1, 2021 (depicted as red dots) excluded from the fitting process. As shown in Figure 5, the observed data from April 1, 2021, fall within the 95% credible interval of the forecasted range, indicating that the model’s predictions align well with the actual trend.
@@
The bottom plot in Figure 5 shows the weekly-averaged effective reproduction number, $R_{\text{eff}}(t)$, with corresponding daily estimates (not averaged) presented in Figure C.1. As anticipated, $R_{\text{eff}}(t)$ fluctuates in response to changes in the epidemic dynamics. The impact of the first lockdown is clearly evident, as it drives $R_{\text{eff}}(t)$ below one, effectively ending the first wave. A gradual increase in $R_{\text{eff}}(t)$ occurred in June 2020, due to the easing of restrictions. $R_{\text{eff}}(t)$ remained above one until mid-October, corresponding to the second wave of the epidemic. The second round of lockdown measures caused $R_{\text{eff}}(t)$ to decline by November 2020, falling below one once more. The emergence of the more transmissible Alpha variant, combined with increased social interactions during the holiday season, resulted in a sharp rise in cases toward the end of 2020. This was followed by a third round of lockdown measures, which once again brought $R_{\text{eff}}(t)$ below one (see the supplementary material of Jaouimaa et al. (2021) for a description of the individual lockdown).
@@
In Figure 6, we present the sequential estimation of the parameters of model (3.2.1) over 365 days of the pandemic. The graphs display the posterior medians of the parameters, accompanied by their respective uncertainty intervals at each time step. The posterior distributions of the parameters obtained at the final time step (see Table C.1) are aligned with the results found in (Cazelles et al., 2021) using PMCMC methods.
Figure D.1 illustrates the application of the O-SMC2 for real-time monitoring, using partial datasets spanning four distinct time periods: February 29, 2020, to May 30, 2020; May 30, 2020, to August 29, 2020; August 29, 2020, to November 28, 2020; and November 28, 2020, to February 27, 2021. At each stage, the algorithm utilizes the results from the previous O-SMC2 run as prior information to compute the joint posterior $p(x_{i:j},\theta\mid y_{i:j})$. This sequential approach enables the efficient integration of new data while maintaining computational efficiency. Moreover, it significantly improves the methodology described in Dureau et al. (2013), which relied on PMCMC and required the model to be rerun from scratch for each new batch of data.
@@ 4 Conclusion
In this paper, we explored the application of an online variant of the Sequential Monte Carlo Squared (O-SMC2) method for real-time inference in stochastic epidemic models. This approach differs from standard SMC2 by integrating a mechanism that approximates the posterior distribution using a fixed window of recent observations, thereby reducing computational demands while maintaining parameter estimation accuracy. This adaptation makes O-SMC2 particularly suitable for dynamic disease modeling, where timely responses are critical. We demonstrated the effectiveness of O-SMC2 in the stochastic SEIR compartment model, validating its performance on both synthetic simulations and real-world COVID-19 data from Ireland. Our results show that O-SMC2, even with a modest window size, effectively tracks epidemic progression and generates robust online estimates of both static and dynamic epidemiological parameters. Moreover, the sequential nature of this method enables the efficient updating of estimates by using results from previous runs as prior information, eliminating the need to rerun the model from scratch as new data batches become available. This feature underscores the practicality of O-SMC2 for real-time epidemiological surveillance, enabling public health authorities to adapt interventions based on the latest information. However, we also acknowledge certain challenges, particularly in the early stages of disease transmission when available data may be sparse and less informative. This can hinder accurate parameter estimation during the onset of an outbreak. To address these limitations, future work will focus on refining our approach by integrating diverse modeling techniques that can better handle early-stage uncertainties and improve prediction accuracy.
@@ Acknowledgements
This publication has emanated from research conducted with the financial support of Taighde Éireann – Research Ireland under Grant number 21/FFP-P/10123.
Appendix A Approximation of the acceptance probability
@@
$p(y_{1:t}|\theta)$ $=\left(\prod_{s=1}^{t-t_{k}-1}p(y_{s}|y_{1:s-1},\theta)\right)\times\left(\prod_{s=t-t_{k}}^{t}p(y_{s}|y_{1:s-1},\theta)\right)$
$=p(y_{1:t-t_{k}-1}|\theta)\times\left(\prod_{s=t-t_{k}}^{t}p(y_{s}|y_{1:s-1},\theta)\right).$ (24)
For $s\geq t-t_{k}$, if we assume that for a sufficiently large $t_{k}$, the older observations $y_{1:t-t_{k}-1}$ have minimal impact on the filtering estimate of the latent state at time $s-1$, i.e., $p(x_{s-1}|y_{1:s-1},\theta)\approx p(x_{s-1}|y_{t-t_{k}:s-1},\theta)$, then, from the result in (Kantas et al., 2015), we have:
[…]
$=p(y_{s}|y_{t-t_{k}:s-1},\theta).$ (25)
@@
$p(y_{1:t}|\theta)\approx p(y_{1:t-t_{k}-1}|\theta)\times p(y_{t-t_{k}:t}|\theta),$ (26)
where
$p(y_{t-t_{k}:t}|\theta)=\prod_{s=t-t_{k}}^{t}p(y_{s}|y_{t-t_{k}:s-1},\theta)\approx\prod_{s=t-t_{k}}^{t}p(y_{s}|y_{1:s-1},\theta).$ (27)
Given the current $\theta$-particle, the new parameter $\theta^{*}$ at time $t$ is drawn from the region where most of the probability mass is located (Chopin, 2002). Under the assumption that the target parameter vector remains constant over time, and provided that the window size $t_{k}$ is large enough to capture the necessary temporal variability in the observations $y_{t-t_{k}:t}$, it is reasonable to assume that the historical likelihoods $p(y_{1:t-t_{k}-1}|\theta)$ and $p(y_{1:t-t_{k}-1}|\theta^{*})$ are approximately equal. Consequently, their contribution to the likelihood ratio becomes negligible. Under this assumption, the full likelihood ratio simplifies to:
$\dfrac{p(y_{1:t}|\theta^{*})}{p(y_{1:t}|\theta)}\approx\dfrac{p(y_{t-t_{k}:t}|\theta^{*})}{p(y_{t-t_{k}:t}|\theta)}.$ (28)
Thus, using the SMC estimate of the incremental likelihood for a fixed number of particles $N_{x}$ (see Equation (8)), we estimate the windowed likelihood $p(y_{t-t_{k}:t}|\theta)$ as:
$\hat{p}_{N_{x}}(y_{t-t_{k}:t}|\theta)=\prod_{s=t-t_{k}}^{t}\left(\frac{1}{N_{x}}\sum_{i=1}^{N_{x}}w_{s}^{i}(x_{s-1}^{i},x_{s}^{i})\right),$ (29)
@@
$\alpha=\min\left\{1,\dfrac{\hat{p}_{N_{x}}(y_{t-t_{k}:t}|\theta^{*})p(\theta^{*})}{\hat{p}_{N_{x}}(y_{t-t_{k}:t}|\theta)p(\theta)}\times\dfrac{q(\theta|\theta^{*})}{q(\theta^{*}|\theta)}\right\}.$ (30)
Appendix B Posterior estimate in the simulation study
Here, we present the posterior estimates of model parameters from two experimental setups using the O-SMC2 algorithm. The experiments test varying window sizes $t_{k}$ to evaluate the computational efficiency and the impact of window length on parameter inference. For each window size, we report the posterior median and 95% credible intervals, comparing them to SMC2 estimates to assess both inference accuracy and computational savings. In our analysis, we define “Relative Efficiency ” as the ratio of the SMC2 computational time to the O-SMC2 computational time, providing a measure of the time reduction achieved by O-SMC2. This ratio quantifies the improvement in computational efficiency for each $t_{k}$ compared to SMC2, with larger ratios indicating greater computational efficiency.
Table B.1: Summary of parameters in Experiment 1. Posterior median and 95% credible intervals using O-SMC2 with varying window sizes $t_{k}$ at the final time step $T=100$. “Relative Efficiency ” denotes the computational time reduction factor compared to SMC2.

### v2

@@ 3.2.2 Data and model fitting
The data we use are daily confirmed case counts extracted from an up-to-date database of the Irish Health Protection Surveillance Centre (HPSC) (https://COVID19ireland-geohive.hub.arcgis.com/). To integrate these observations into our model, we define the variable: $C^{I}_{t}$ representing the cumulative number of infectious cases. This variable evolves according to the equation:
@@
Hence, $\mu_{t}=C^{I}_{t+1}-C^{I}_{t}$ represents the daily expected number of cases. Since daily reported case counts are subject to observation error and typically exhibit overdispersion (where the conditional variance exceeds the conditional mean), we model the observed daily cases, $Z^{I}_{t}$, using a normal approximation of the negative binomial distribution (chosen for computational efficiency, Funk et al. (2018)). The negative binomial distribution is a generalization of the Poisson distribution that can account for overdispersion in count data. Specifically, for the negative binomial distribution, the mean is $\mu_{t}$ and the variance is $\mu_{t}(1+\phi\mu_{t})$, where $\phi$ is an overdispersion parameter. To avoid singularities when $\mu_{t}=0$, we followed the approach of Funk et al. (2018), rounding variances smaller than 1 up to 1. The observed daily case counts follow a negative binomial distribution:
@@
We estimate the spread of COVID-19 in Ireland from February 29th, 2020, to February 28th, 2021, using the daily new case count data over 365 days. The Irish population in 2020 was estimated to be $N=4,965,439$. The O-SMC2 uses $N_{\theta}=1000$ parameter particles, $N_{x}=500$ state particles, $M=5$ successive PMMH moves and a window size of $t_{k}=80$ days. Table 1 in Appendix D gives the description and prior of the inferred parameters used in the model (3.2.1). These priors are based on previous studies of COVID-19 dynamics in Ireland (Irish Department of Health, 2020; Cazelles et al., 2021). For the choice of prior distributions for the hyperparameters, $\nu_{\beta}$ (which governs the variability in the transmission rate) and $\phi$ (the overdispersion parameter), we adopt settings that allow for rapid changes in the transmission rate, ensuring the model can adapt to sudden shifts in disease dynamics, as discussed in (Funk et al., 2018). Regarding initial states, we set: $S_{0}=N-E_{0}-A_{0}$, $E_{0}=1$, $A_{0}\sim\mathcal{U}(\{10,\dots,50\})$ and $I_{0}=R_{0}=0$. The execution time for the 365 days considered here was less than 5 hours, making it suitable for practical real-time applications.
@@
Figure 6 illustrates the model’s fit to the daily number of COVID-19 cases reported in Ireland. The top plot shows the daily new cases, with three noticeable peaks: one in mid-April 2020, another in mid-November 2020, and a larger peak in early January 2021. The model successfully captures these three waves, with observed data (orange dots) falling within the 95% credible interval on most days. The fitting period achieved a mean absolute error (MAE) of 139.59 and a continuous ranked probability score (CRPS) of 105.78. These evaluation metrics are defined in detail in Appendix D. However, a sudden drop in reported cases around mid-January 2021 is not fully captured, likely due to real-world factors such as reporting delays or testing changes not included in the model. Beyond February 28th, 2021, the model projects cases forward for 21 days, excluding data from April 1 (red dots) from the fitting process. The forecast period shows good predictive accuracy with an MAE of 113.71 and a CRPS of 98.45, and the observed data fall within the 95% credible interval, indicating strong agreement between predictions and actual trends.
@@
The bottom plot in Figure 6 shows the weekly-averaged effective reproduction number, $R_{\text{eff}}(t)$, with corresponding daily estimates (not averaged) presented in Figure 1 in Appendix D. As anticipated, $R_{\text{eff}}(t)$ fluctuates in response to changes in the epidemic dynamics. The impact of the first lockdown is clearly evident, as it drives $R_{\text{eff}}(t)$ below one, effectively ending the first wave. A gradual increase in $R_{\text{eff}}(t)$ occurred in June 2020, due to the easing of restrictions. $R_{\text{eff}}(t)$ remained above one until mid-October, corresponding to the second wave of the epidemic. The second round of lockdown measures caused $R_{\text{eff}}(t)$ to decline by November 2020, falling below one once more. The emergence of the more transmissible Alpha variant, combined with increased social interactions during the holiday season, resulted in a sharp rise in cases toward the end of 2020. This was followed by a third round of lockdown measures, which once again brought $R_{\text{eff}}(t)$ below one (see the supplementary material of Jaouimaa et al. (2021) for a description of the individual lockdown measures).
@@
In Figure 7, we present the sequential estimation of the parameters of model (3.2.1) over 365 days of the pandemic. The graphs display the posterior medians of the parameters, accompanied by their respective uncertainty intervals at each time step. The posterior distributions of the parameters obtained at the final time step (see Table 1 in Appendix D) are aligned with the results found in Cazelles et al. (2021) using PMCMC methods.
@@
Figure 1 in Appendix E illustrates the application of the O-SMC2 for real-time monitoring, using partial datasets spanning four distinct time periods: February 29, 2020, to May 30, 2020; May 30, 2020, to August 29, 2020; August 29, 2020, to November 28, 2020; and November 28, 2020, to February 27, 2021. At each stage, the algorithm utilizes the results from the previous O-SMC2 run as prior information to compute the joint posterior $p(x_{i:j},\theta\mid y_{i:j})$. This sequential approach enables the efficient integration of new data while maintaining computational efficiency. Moreover, it significantly improves the methodology described in Dureau et al. (2013), which relied on PMCMC and required the model to be rerun from scratch for each new batch of data.
@@ 4 Conclusion
In this paper, we explored the application of an online variant of the Sequential Monte Carlo Squared (O-SMC2) method for real-time inference in stochastic epidemic models. This approach differs from standard SMC2 by integrating a mechanism that approximates the posterior distribution using a fixed window of recent observations, thereby reducing computational demands while maintaining parameter estimation accuracy. This adaptation makes O-SMC2 particularly suitable for dynamic disease modeling, where timely responses are critical. We demonstrated the effectiveness of O-SMC2 in the stochastic SEIR compartment model, validating its performance on both synthetic simulations and real-world COVID-19 data from Ireland. Our results show that O-SMC2, even with a modest window size, effectively tracks epidemic progression and generates robust online estimates of both static and dynamic epidemiological parameters. Moreover, the sequential nature of this method enables the efficient updating of estimates by using results from previous runs as prior information, eliminating the need to rerun the model from scratch as new data becomes available incrementally. This feature underscores the practicality of O-SMC2 for real-time epidemiological surveillance, informing public health authorities on adaptation of interventions based on real-time information assimilation. However, we also acknowledge certain challenges, particularly in the early stages of an outbreak when available data may be sparse and less informative. During this phase, prior uncertainty, potential parameter non-identifiability, and low case counts can collectively reduce the reliability of estimates. These issues can result in wide credible intervals and greater variability in posterior distributions. One practical consideration is the choice of the fixed window size, $t_{k}$, which impacts the balance between computational efficiency and the informativeness of the data used for inference. In our experiments, setting $t_{k}$ to approximately 20% of the total time series length provided a reasonable compromise. However, in truly online settings, where data arrives as a continuous stream without a predetermined end, a fixed window size may not be optimal. An adaptive strategy that dynamically adjusts $t_{k}$, for example, by increasing the window size when the PMMH acceptance rate fall below a certain threshold, could improve the robustness during periods of rapid change or less informative observations. Exploring such adaptive mechanisms and integrating more flexible modeling approaches are key directions for enhancing the responsiveness and reliability of early-stage inference in real-time epidemic monitoring.
@@ Acknowledgements
This publication has emanated from research conducted with the financial support of Taighde Éireann – Research Ireland under Grant number 21/FFP-P/10123. We wish to thank the editor and the anonymous reviewers for their constructive comments and suggestions which greatly improved the quality of this paper.
Appendix A Structural identifiability analysis
To assess structural identifiability, we first derived a deterministic mean-field approximation of the stochastic SEIR model. This step is necessary because most identifiability analysis tools require a system of ordinary differential equations (ODEs) as input. Starting from the discrete-time stochastic model, we replaced each binomial transition with its expected value and applied the first-order approximation $1-e^{-x}\approx x$ for $x\ll 1$. Taking the limit as $\delta t\to 0,$ , we obtained the following system:
$\frac{dS}{dt}=-\beta\frac{SI}{N},\quad\frac{dE}{dt}=\beta\frac{SI}{N}-\sigma E,\quad\frac{dI}{dt}=\sigma E-\gamma I,\quad\frac{dR}{dt}=\gamma I.$ (24)
We assumed the observable is $y_{1}(t)=\sigma E(t)$, representing daily new infections. All parameters are also assumed to be static throughout the analysis, consistent with standard structural identifiability frameworks.
Structural identifiability refers to the theoretical ability to uniquely recover model parameters from perfect data, given the model structure and observables. This was assessed using the Julia package StructuralIdentifiability.jl (Dong et al., 2023). The total population size $N$ and the initial conditions were assumed to be known during the analysis. Results (Table 1) show that for both the standard SEIR and the SEAIR model, key parameters such as $\beta$, $\sigma$, $\gamma$ and additional parameters in the extended model, are structurally globally identifiable. For the SEIR model, we also found that all parameters remain globally identifiable even when the initial conditions are unknown, provided that N is known. These findings confirm that the selected observables are theoretically sufficient to uniquely identify the parameters prior to applying the O-SMC2 algorithm. For a tutorial introduction to StructuralIdentifiability.jl in epidemic modeling, see (Liyanage et al., 2025).
Table 1: Structural identifiability results.
SEIR SEIAR
julia> ode = @ODEmodel(
       S’(t) = - beta*S(t)*I(t)/N,
       E’(t) = beta*S(t)*I(t)/N - sigma*E(t),
       I’(t) = sigma*E(t) - gamma*I(t),
       R’(t) = gamma*I(t),
       y1(t) = sigma*E(t),
       y2(t) = N
julia> assess_identifiability(ode, known_ic = [S,E,I,R])
  S(0)  => :globally
  E(0)  => :globally
  I(0)  => :globally
  R(0)  => :globally
  beta  => :globally
  gamma => :globally
  sigma => :globally
julia> ode = @ODEmodel(
       S’(t) = - beta*S(t)*(I(t)+ra*A(t))/N,
       E’(t) = beta*S(t)*(I(t)+ra*A(t))/N - sigma*E(t),
       A’(t) = pa*sigma*E(t) - gamma*A(t),
       I’(t) = (1-pa)*sigma*E(t) - gamma*I(t),
       R’(t) = gamma*A(t) + gamma*I(t),
       y1(t) = (1-pa)*sigma*E(t),
       y2(t) = N
julia> assess_identifiability(ode, known_ic = [S,E,A,I,R])
  S(0)  => :globally
  E(0)  => :globally
  A(0)  => :globally
  I(0)  => :globally
  R(0)  => :globally
  N     => :globally
  beta  => :globally
  gamma => :globally
  pa    => :globally
  ra    => :globally
  sigma => :globally
Appendix B Approximation of the acceptance probability
@@
$p(y_{1:t}|\theta)$ $=\left(\prod_{s=1}^{t-t_{k}}p(y_{s}|y_{1:s-1},\theta)\right)\times\left(\prod_{s=t-t_{k}+1}^{t}p(y_{s}|y_{1:s-1},\theta)\right)$
$=p(y_{1:t-t_{k}}|\theta)\times\left(\prod_{s=t-t_{k}+1}^{t}p(y_{s}|y_{1:s-1},\theta)\right).$ (25)
For $s\geq t-t_{k}$, if we assume that for a sufficiently large $t_{k}$, the older observations $y_{1:t-t_{k}}$ have minimal impact on the filtering estimate of the latent state at time s1s-1, i.e., $p(x_{s-1}|y_{1:s-1},\theta)\approx p(x_{s-1}|y_{t-t_{k}+1:s-1},\theta)$, then, from the result in (Kantas et al., 2015), we have:
$p(y_{s}|y_{1:s-1},\theta)$ $=\int g(y_{s}|x_{s},\theta)f(x_{s}|x_{s-1},\theta)p(x_{s-1}|y_{1:s-1},\theta)\,dx_{s-1:s}$
$\approx\int g(y_{s}|x_{s},\theta)f(x_{s}|x_{s-1},\theta)p(x_{s-1}|y_{t-t_{k}+1:s-1},\theta)\,dx_{s-1:s}$
$=p(y_{s}|y_{t-t_{k}+1:s-1},\theta).$ (26)
@@
$p(y_{1:t}|\theta)\approx p(y_{1:t-t_{k}}|\theta)\times p(y_{t-t_{k}+1:t}|\theta),$ (27)
with:
$p(y_{t-t_{k}+1:t}|\theta)=\prod_{s=t-t_{k}+1}^{t}p(y_{s}|y_{t-t_{k}+1:s-1},\theta)\approx\prod_{s=t-t_{k}+1}^{t}p(y_{s}|y_{1:s-1},\theta).$ (28)
Given the current $\theta$-particle, the new parameter $\theta^{*}$ at time $t$ is drawn from the region where most of the probability mass is located (Chopin, 2002). Under the assumption that the target parameter vector remains constant over time, and provided that the window size $t_{k}$ is large enough to capture the necessary temporal variability in the observations $y_{t-t_{k}+1:t}$, it is reasonable to assume that the historical likelihoods $p(y_{1:t-t_{k}}|\theta)$ and $p(y_{1:t-t_{k}}|\theta^{*})$ are approximately equal. Consequently, their contribution to the likelihood ratio becomes negligible. Under this assumption, the full likelihood ratio simplifies to:
$\dfrac{p(y_{1:t}|\theta^{*})}{p(y_{1:t}|\theta)}\approx\dfrac{p(y_{t-t_{k}+1:t}|\theta^{*})}{p(y_{t-t_{k}+1:t}|\theta)}.$ (29)
Thus, using the SMC estimate of the incremental likelihood for a fixed number of particles $N_{x}$ (see Equation (8)), we estimate the windowed likelihood $p(y_{t-t_{k}+1:t}|\theta)$ as:
$\hat{p}_{N_{x}}(y_{t-t_{k}+1:t}|\theta)=\prod_{s=t-t_{k}+1}^{t}\left(\frac{1}{N_{x}}\sum_{i=1}^{N_{x}}w_{s}^{i}(x_{s-1}^{i},x_{s}^{i})\right),$ (30)
@@
$\alpha=\min\left\{1,\dfrac{\hat{p}_{N_{x}}(y_{t-t_{k}+1:t}|\theta^{*})p(\theta^{*})}{\hat{p}_{N_{x}}(y_{t-t_{k}+1:t}|\theta)p(\theta)}\times\dfrac{q(\theta|\theta^{*})}{q(\theta^{*}|\theta)}\right\}.$ (31)
Appendix C Posterior estimate in the simulation study
Here, we present the posterior estimates of model parameters from two experimental setups using the O-SMC2 algorithm. The experiments test varying window sizes $t_{k}$ to evaluate the computational efficiency and the impact of window length on parameter inference. Table 1 summarizes the parameter settings and prior distributions used in the two simulation experiments. For each window size, we report the posterior median and 95% credible intervals, comparing them to SMC2 estimates to assess both inference accuracy and computational savings. Tables 2 and 3 display these results for Experiments 1 and 2, respectively, including the “Relative Efficiency” defined as the ratio of the SMC2 computational time to the O-SMC2 computational time. This metric quantifies the time reduction achieved by O-SMC2 for each $t_{k}$ compared to SMC2, with larger ratios indicating greater computational efficiency.
Table 1: Summary of parameter settings used in the two simulation Experiments. $\mathcal{U}$(inf, sup) indicates a uniform distribution and $\mathcal{TN}$ stands for truncated normal distribution ($\mathcal{TN}_{[\text{inf, sup}]}(\text{mean, std})$).
Parameter Description Value in simulation Prior Notes
Experiment 1
$\beta$ Transmission rate $0.6$ $\mathcal{U}(0,1)$ Estimated
$\sigma$ Rate from exposed to infectious $1/3$ $\mathcal{U}(0,1)$ Estimated
$\gamma$ Recovery rate $1/5$ $\mathcal{U}(0,1)$ Estimated
$I_{0}$ Initial infected individuals 1 $\mathcal{U}(\{0,\dots,5\})$
NN Population size 60006000 Fixed
Experiment 2
$\beta(t)$ Time-varying transmission rate $\beta(t)=0.3\exp\left(\sin\left(\frac{2\pi t}{55}\right)-\frac{t}{80}\right)$ Modeled via Eq. (16) Estimated
$\sigma$ Rate from exposed to infectious $1/2$ $\mathcal{U}(0.3,0.6)$ Estimated
$\gamma$ Recovery rate $1/7$ $\mathcal{TN}_{[0,1]}(0.14,0.01)$ Estimated
$\nu_{\beta}$ Volatility in $\beta(t)$ $\mathcal{TN}_{[0.05,0.15]}(0.1,0.05)$ Estimated
$\beta(0)$ Initial transmission rate 0.3 $\mathcal{U}(0.2,0.5)$
$I_{0}$ Initial infected individuals 55 $\mathcal{U}(\{0,\dots,5\})$
NN Population size $200,000$ Fixed
Table 2: Summary of parameter estimates in Experiment 1. Posterior medians and 95% credible intervals were obtained using O-SMC2 with varying window sizes $t_{k}$ at the final time step $T=100$. “Relative Efficiency” denotes the factor by which computational time is reduced compared to SMC2.